OpenStudy (anonymous):

substitution 2. 3r+t=5, 2r-t=5

OpenStudy (cgreenwade2000):

Use the first equation to solve for t.

jimthompson5910 (jim_thompson5910):

3r+t=5 t=5-3r ------------- 2r-t=5 2r-(5-3r)=5 2r - 5 + 3r = 5 Keep going to solve for r

OpenStudy (anonymous):

thx jim

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

once you have the solution for r, use it to find the solution for t

OpenStudy (anonymous):

ok thx